OpenStudy (anonymous):

write a linear function f such that f (-8)=3 and is a) parallel to the line 2x+3y=5 b) perpendicular to the line 2x+3y=5

OpenStudy (anonymous):

a) the given line is 2x +3y =5 the given point is (-8,3) now, plug the point into the left hand side of the line, you get 2*(-8) + 3*(3) = -7 so that 2x + 3y =-7 is the required function b) to this problem, you must find out the slope of the given line, then take its reciprocal to get the slope of the new one. let say it is m' with m' and the given point (-8,3) you can write out the function